I am in algebra 2 and we are learning about "interperetting graphs. Equations like, "If f(x)=Xsquared the the graph of each equatrion below is a parabola. Describe the location of the parabola relative to the graph of f(x)=xsquared.... y=f(x)-3 or y=f(x-2)" im so lost, we have a sustitute because our teacher is on meternity leave on he knows nothing about anything he is teaching. Please help!
Answer by scott8148(6628)

one form of the equation for a parabola is y=a(x-h)^2+k

this is a parabola with the vertex (nose) at (h,k)

changing the h value moves the parabola horizontally; and changing the k value moves the parabola vertically

the a value determines the width (fat or pointy) of the parabola and the direction of the opening (up or down)

so y=x^2 is parabola with the vertex at the origin (0,0)opening upward

y=(x-2)^2 (y=x^2-4x+4) is a parabola with the vertex at (2,0)

y=(x-2)^2+3 (y=x^2-4x+7) has the vertex at (2,3)